a teacher chooses a student at randmon from a class of 10 boys and 15 girls.Find the odd in favor of,and the odd against,each event.
(a)choosing a girl
(b)choosing a boy

Answer by stanbon(75887)

a teacher chooses a student at random from a class of 10 boys and 15 girls.Find the odds in favor of,and the odds against,each event.
(a)choosing a girl
P(choosing girl) = 15/25 = 3/5
P(not choosing girl) = 2/5
Odds in favor of (3/5)/(2/5) = 3:2
Odds against: (2/5)/(3/5) = 2:3
---------------------------------
(b)choosing a boy
Prob(choosing boy) = 10/25 = 2/5
Prob(not choosing boy) = 3/5
--------------
Odds in favor of = (2/5)/(3/5) = 2:3
Odds against: (3/5)/(2/5) = 3/2
===============================
Cheers,
Stan H.